OpenStudy (anonymous):

Janet said that the common difference in the sequence below is -2. Is she correct? Justify your answer using complete sentences and proper grammar. 1, -2, 4, -8....

OpenStudy (freckles):

well what are the following equal to term-previous term: -2-1= 4-(-2)= -8-4=

OpenStudy (freckles):

if they are all the same number then the sequence does have a common difference and whatever number that is is the common difference

OpenStudy (anonymous):

So yeah

OpenStudy (freckles):

did you actually perform the differences above?

OpenStudy (anonymous):

yes .

OpenStudy (freckles):

-2-1=? 4-(-2)=? -8-4=?

OpenStudy (freckles):

so you got each one of those is -2?

OpenStudy (anonymous):

No

OpenStudy (freckles):

right the answer is no the sequence doesn't have a common difference because -2-1 is not the same as 4-(-2) or the same as -8-4

OpenStudy (anonymous):

but i thought because when multiplied by -2 idk i thought multiplying made sense

OpenStudy (freckles):

common difference assumes you have an arithemetic sequence and we don't

OpenStudy (anonymous):

ooohhhhhhhhh

OpenStudy (freckles):

common ratio assumes we have a geometric sequence which we do

OpenStudy (freckles):

the word that makes the whole thing false is the difference word

OpenStudy (freckles):

if they said common ratio then it would have been true

OpenStudy (anonymous):

ohh , thank you so much .

OpenStudy (freckles):

since term/previous term =-2 that is -2/1=4/-2=-8/4=-2

OpenStudy (freckles):

and np